# Optimization of the Downlink Spectral- and Energy-Efficiency of RIS-aided Multi-user URLLC MIMO Systems

### v1

@@ Abstract
Modern wireless communication systems are expected to provide improved latency and reliability. To meet these expectations, a short packet length is needed, which makes the first-order Shannon rate an inaccurate performance metric for such communication systems. A more accurate approximation of the achievable rates of finite-block-length (FBL) coding regimes is known as the normal approximation (NA). It is therefore of substantial interest to study the optimization of the FBL rate in multi-user multiple-input multiple-output (MIMO) systems, in which each user may transmit and/or receive multiple data streams. Hence, we formulate a general optimization problem for improving the spectral and energy efficiency of multi-user MIMO-aided ultra-reliable low-latency communication (URLLC) systems, which are assisted by reconfigurable intelligent surfaces (RISs). We show that a RIS is capable of substantially improving the performance of multi-user MIMO-aided URLLC systems. Moreover, the benefits of RIS increase as the packet length and/or the tolerable bit error rate are reduced. This reveals that RISs can be even more beneficial in URLLC systems for improving the FBL rates than in conventional systems approaching Shannon rates.
@@ I Introducton
The sixth generation (6G) of wireless communication systems is expected to significantly improve the spectral efficiency (SE), energy efficiency (EE), and reliability of the existing systems, despite of providing a lower latency than 5G [1, 2]. Thus, 6G should employ radical new technologies such as reconfigurable intelligent surface (RIS) to meet these expectations. Moreover, 6G has to support a large variety of applications, which require ultra-reliable and low-latency communications (URLLC) [1, 2]. To attain low latency, realistic finite block length (FBL) codes have to be employed. In this content, the classic Shannon rate is an inaccurate performance metric for URLLC systems. Indeed, the FBL rate is more challenging to optimize than the Shannon rate, especially in multiple-input multiple-output (MIMO) systems, when multi-stream communication is targeted. In fact, to the best of our knowledge, resource allocation has not been designed for multi-user MIMO (MU-MIMO) systems relying on FBL coding in the open literature for the scenario of multiple streams per user. Developing suitable resource allocation schemes is even more challenging in RIS-assisted systems since this requires the joint optimization of the transmit covariance matrices and the channels, which depend on the RIS elements. To close this knowledge gap, we derive a closed-form expression for the rates of users in MU-MIMO systems using realistic FBL coding when multiple streams are allowed. Then, we develop an optimization framework for MU-MIMO RIS-assisted URLLC systems and show that a RIS can substantially improve the SE and EE. Our results show that a RIS can be even more beneficial in MIMO URLLC systems than in systems approaching the classic Shannon rate, since a RIS provides higher gains for short packets and/or for low tolerable bit error rates (BERs).
@@ I-A Literature Review
A main goal of 6G is to drastically enhance the SE and EE, which are even more vital for applications related to FBL coding. To realize this goal, 6G has to employ powerful emerging technologies such as RISs as well as existing MIMO solutions [3, 4]. A RIS was shown to be able to substantially enhance the EE and SE [5, 6, 7, 8, 9, 10], when studying different performance metrics such as the sum rate, minimum rate, total power consumption required for achieving a specific quality of service (QoS), minimum signal-to-interference-plus-noise ratio (SINR), global EE (GEE), and interference leakage. For instance in [5], the authors proposed algorithms for increasing the GEE and sum rate of a multiple-input single-output (MISO) broadcast channel (BC). In [6], it was shown that a RIS reduces the power consumption of the single-cell MISO BC when a minimum SINR per user has to be ensured. Moreover, the authors of [7] showed that a RIS is capable of increasing the minimum SINR of a single-cell MISO BC for a given power budget. The benefit of RIS in terms of both the minimum rate and the minimum EE was studied in [8] for a multi-cell MISO BC employing non-orthogonal multiple access (NOMA). Furthermore, the authors of [9] studied the interference leakage in a MIMO-aided $K$-user interference channel assisted by RIS, showing that an appropriately designed RIS mitigates the total interference power of the system.
@@
The performance of RIS relying on the Shannon capacity achieving codes has been studied in [5, 6, 7, 8, 9], but naturally, a RIS can also be beneficial in multi-user systems using FBL coding, as shown in [11, 12, 13, 14, 15, 16]. For instance, in [11], resource allocation schemes were developed for MISO URLLC systems, assisted by simultaneously transmit and reflect (STAR-) RIS, and it was shown that a RIS (either STAR or purely reflective) substantially improves the SE and EE of URLLC systems. In [12], it was demonstrated that RIS and rate splitting can be mutually beneficial tools of enhancing the EE and SE performance of interference-limited URLLC systems. In [14], the advantage of employing a RIS and NOMA in a two-user single-input single-output (SISO) URLLC BC was shown.
@@
All the aforementioned papers [11, 12, 13, 14, 15, 16] studied multi-user RIS-assisted URLLC systems, but only supported single-stream data transmission per user. However, RIS can also improve the system performance when parallel frequency-domain channels are employed along with FBL coding [17]. Nevertheless, resource allocation has not been designed for multi-user MIMO URLLC systems supporting multiple streams per user in the open literature. This is particularly challenging for RIS-assisted systems. Indeed, only a few treatises exist in multiple-stream data transmission in MIMO systems, which mainly studied a single-user scenario without considering RISs [18, 19, 20]. Thus, multi-user MIMO systems both with and without RISs require further investigations. MIMO systems support multiple-stream data transmissions per user, which exploit the spectrum efficiently and improve the EE at a specific QoS. Below, we briefly describe the challenges of optimizing the FBL rates when multiple-stream data transmission is supported.
@@
In the FBL regime, the achievable rate depends not only on the Shannon rate, $C$, but also on the channel’s dispersion, $V$, the packet length, $n_{t}$, and the tolerable bit error rate $\epsilon$. An accurate approximation for the achievable rate of FBL coding in parallel channels is the normal approximation (NA) [^1] Note that the NA may not be accurate when the packet length is extremely short, and/or the tolerable bit error rate is extremely low. For further discussions regarding the accuracy of the NA, please refer to [21, 22, 23]., which is given by [24, Theorem 78]
@@
Multi-user Channel dispersion in [27] EE MIMO Multiple streams STAR-RIS
This paper $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
[12] $\surd$ $\surd$ $\surd$ - - -
[28] $\surd$ $\surd$ - - - -
[13, 14, 15, 16, 17, 29] $\surd$ - - - - -
[30, 31] - - - - - -
@@ I-C Contribution
To elaborate, our optimization framework is flexible and can be utilized in any MU-MIMO URLLC system aided by RISs. Furthermore, the framework may be used for solving a broad spectrum of optimization problems for which the objective function and/or constraints are linear functions of the rates and/or EE of users. The convergence of our framework is ensured towards a stationary point for the general optimization problem, when the feasibility set for the RIS elements conforms to a convex set. We consider a multi-cell MIMO BC as an example of the networks that our framework can be applied to. Furthermore, we consider both the EE and SE metrics for investigating the performance of RIS in MU-MIMO URLLC systems. For the SE metric, we consider the sum rate and the minimum rate of the users, which are among the most common performance metrics for SE. Moreover, we evaluate the EE by optimizing the GEE and the minimum EE of users. Note that the sum rate and global EE are pivotal overall system performance metrics. By contrast, the minimum rate and the EE of specific users consider the individual performance of the users and can provide reasonable rate/EE-fairness among the users since typically all the users are allocated similar rate/EE when the minimum rate/EE is maximized. Thus, considering all these metrics can provide a complete picture of the performance of MU-MIMO URLLC systems aided by RISs. Moreover, we make realistic assumptions regarding the channel models and the feasible sets of the RIS coefficients for appropriately examining the RIS performance.
@@ II System model
Our proposed framework can be applied to any RIS-assisted MU-MIMO URLLC system that treats interference as noise at the receivers. As an example of such MU-MIMO systems, we consider a multicell MIMO RIS-assisted downlink (DL) BC comprising $L$ BSs, as shown in Fig. 1. We assume that each BS has $N_{BS}$ the DL transmission antennas (TAs) and serves $K$ users having $N_{u}$ receive antennas (RAs) each. Furthermore, we assume that there are $M$ reflective passive RISs to assist the BSs, and each RIS has $N_{RIS}$ elements. Additionally, we assume having perfect, instantaneous, and global CSI. For notational simplicity, we consider a symmetric scenario in which all the BSs/users/RISs employ similar devices without loss of generality. However, the framework can be applied to asymmetric systems having an arbitrary number of antennas at the transmit and receive sides, and different number of users per cell.
Figure 1: A multicell BC assisted by RISs.
@@ II-A RIS Model
[…]
$\mathbf{H}_{lk,i}\!\left(\{\bf{\Psi}\}\right)\!\!=\!\!\underbrace{\sum_{m=1}^{M}\mathbf{G}_{lk,m}{\bf\Psi}_{m}\mathbf{G}_{m,i}}_{\text{Link through RIS}}+\!\!\!\underbrace{\mathbf{F}_{lk,i}}_{\text{Direct link}}\!\!\!\!\in\mathbb{C}^{N_{u}\times N_{BS}}\!\!,$ (2)
[…]
${\bf\Psi}_{m}=\text{diag}\left(\psi_{m_{1}},\psi_{m_{2}},\cdots,\psi_{m_{N_{RIS}}}\right).$
Assuming having passive RISs, the absolute value of the RIS coefficients cannot be greater than 1, which results in the following set for the feasible RIS coefficients [3, Eq. (11)]
@@
In this feasibility set, the amplitude and phase of each RIS element are assumed to be independent optimization variables, which might not be realistic. Another common assumption is that the RIS coefficients have to adhere to the unit modulus constraint [4, 3, 6, 7, 33, 32, 34], which leads to the set
@@
[…]
@@
where we have ${\bf\Psi}_{m}^{r}=\text{diag}\left(\psi_{m_{1}}^{r},\psi_{m_{2}}^{r},\cdots,\psi_{m_{N_{RIS}}^{r}}\right)$ and ${\bf\Psi}_{m}^{t}=\text{diag}\left(\psi_{m_{1}}^{t},\psi_{m_{2}}^{t},\cdots,\psi_{m_{N_{RIS}}^{t}}\right)$. Assuming operating in a passive mode, the absolute values of the reflection and transmission coefficients have to satisfy
$|\psi_{m_{n}}^{r}|^{2}+|\psi_{m_{n}}^{t}|^{2}\leq 1,\hskip 28.45274pt\forall mn,$ (6)
@@
There are three different STAR-RIS schemes, including the energy splitting (ES), mode switching (MS) and time switching (TS) schemes [37, 36]. Since the main focus of this work is on evaluating the impact of employing multiple streams per user regimes, we consider only the MS scheme. Note that the MS scheme has a lower implementation complexity than the ES scheme, but its performance may be comparable to that of the ES scheme as shown in, e.g., [38, 39, 40]. The framework proposed in this treatise can be extended to include the ES and TS schemes by following an approach similar to [39, 40]. Note that there are other RIS technologies and/or more practical feasibility sets for STAR/reflective RIS as mentioned in [3, 41], which should be considered in future studies.
It is worth emphasizing that the channel matrices are assumed to be linear/affine functions of the RIS elements. To simplify the notations/equations, we remove this dependency and subsequently denote the channels as $\mathbf{H}_{lk,i}$ for all $l,k,i$, hereafter. Additionally, we denote the set of the feasible RIS elements as $\mathcal{T}$, unless we refer to a specific set.

### v2

@@ Abstract
Modern wireless communication systems are expected to provide improved latency and reliability. To meet these expectations, a short packet length is needed, which makes the first-order Shannon rate an inaccurate performance metric for such communication systems. A more accurate approximation of the achievable rates of finite-block-length (FBL) coding regimes is known as the normal approximation (NA). It is therefore of substantial interest to study the optimization of the FBL rate in multi-user multiple-input multiple-output (MIMO) systems, in which each user may transmit and/or receive multiple data streams. Hence, we formulate a general optimization problem for improving the spectral and energy efficiency of multi-user MIMO-aided ultra-reliable low-latency communication (URLLC) systems, which are assisted by reconfigurable intelligent surfaces (RISs). We show that an RIS is capable of substantially improving the performance of multi-user MIMO-aided URLLC systems. Moreover, the benefits of RIS increase as the packet length and/or the tolerable bit error rate are reduced. This reveals that RISs can be even more beneficial in URLLC systems for improving the FBL rates than in conventional systems approaching Shannon rates.
@@ I Introducton
The sixth generation (6G) of wireless communication systems is expected to significantly improve the spectral efficiency (SE), energy efficiency (EE), and reliability of the existing systems, despite of providing a lower latency than 5G [1, 2]. Thus, 6G should employ radical new technologies such as reconfigurable intelligent surface (RIS) to meet these expectations. Moreover, 6G has to support a large variety of applications, which require ultra-reliable and low-latency communications (URLLC) [1, 2]. To attain low latency, realistic finite block length (FBL) codes have to be employed. In this content, the classic Shannon rate is an inaccurate performance metric for URLLC systems. Indeed, the FBL rate is more challenging to optimize than the Shannon rate, especially in multiple-input multiple-output (MIMO) systems, when multi-stream communication is targeted. In fact, to the best of our knowledge, resource allocation has not been designed for multi-user MIMO (MU-MIMO) systems relying on FBL coding in the open literature for the scenario of multiple streams per user. Developing suitable resource allocation schemes is even more challenging in RIS-assisted systems since this requires the joint optimization of the transmit covariance matrices and the channels, which depend on the RIS elements. To close this knowledge gap, we derive a closed-form expression for the rates of users in MU-MIMO systems using realistic FBL coding when multiple streams are allowed. Then, we develop an optimization framework for MU-MIMO RIS-assisted URLLC systems and show that an RIS can substantially improve the SE and EE. Our results show that an RIS can be even more beneficial in MIMO URLLC systems than in systems approaching the classic Shannon rate, since an RIS provides higher gains for short packets and/or for low tolerable bit error rates (BERs).
@@ I-A Literature Review
A main goal of 6G is to drastically enhance the SE and EE, which are even more vital for applications related to FBL coding. To realize this goal, 6G has to employ powerful emerging technologies such as RISs as well as existing MIMO solutions [3, 4]. An RIS was shown to be able to substantially enhance the EE and SE [5, 6, 7, 8, 9, 10], when studying different performance metrics such as the sum rate, minimum rate, total power consumption required for achieving a specific quality of service (QoS), minimum signal-to-interference-plus-noise ratio (SINR), global EE (GEE), and interference leakage. For instance in [5], the authors proposed algorithms for increasing the GEE and sum rate of a multiple-input single-output (MISO) broadcast channel (BC). In [6], it was shown that an RIS reduces the power consumption of the single-cell MISO BC when a minimum SINR per user has to be ensured. Moreover, the authors of [7] showed that an RIS is capable of increasing the minimum SINR of a single-cell MISO BC for a given power budget. The benefit of RIS in terms of both the minimum rate and the minimum EE was studied in [8] for a multi-cell MISO BC employing non-orthogonal multiple access (NOMA). In a multi-cell BC, each user is associated with only a single base station (BS) at a time. In [11, 12, 13], it was shown that RISs can also improve the performance of cell-free systems in which each user simultaneously communicates with several access points, rather than a single BS.
@@
The performance of RIS relying on the Shannon capacity achieving codes has been studied in [5, 6, 7, 8, 9], but naturally, an RIS can also be beneficial in multi-user systems using FBL coding, as shown in [14, 15, 16, 17, 18, 19]. For instance, in [14], resource allocation schemes were developed for MISO URLLC systems, assisted by simultaneously transmit and reflect (STAR-) RIS, and it was shown that an RIS (either STAR or purely reflective) substantially improves the SE and EE of URLLC systems. In [15], it was demonstrated that RIS and rate splitting can be mutually beneficial tools of enhancing the EE and SE performance of interference-limited URLLC systems. In [17], the advantage of employing an RIS and NOMA in a two-user single-input single-output (SISO) URLLC BC was shown.
@@
The papers [14, 15, 16, 17, 18, 19] studied multi-user RIS-assisted URLLC systems, but only supported single-stream data transmission per user. However, RIS can also improve the system performance when parallel frequency-domain channels are employed along with FBL coding [20]. Nevertheless, resource allocation has not been designed for multi-user MIMO URLLC systems supporting multiple streams per user in the open literature. This is particularly challenging for RIS-assisted systems. Indeed, only a few treatises exist in multiple-stream data transmission in MIMO systems, which mainly studied a single-user scenario without considering RISs [21, 22, 23]. Thus, multi-user MIMO systems both with and without RISs require further investigations. MIMO systems support multiple-stream data transmissions per user, which exploit the spectrum efficiently and improve the EE at a specific QoS. Below, we briefly describe the challenges of optimizing the FBL rates when multiple-stream data transmission is supported.
@@
In the FBL regime, the achievable rate depends not only on the Shannon rate, $C$, but also on the channel’s dispersion, $V$, the packet length, $n_{t}$, and the tolerable bit error rate $\epsilon$. An accurate approximation for the achievable rate of FBL coding in parallel channels is the normal approximation (NA) [^1] Note that the NA may not be accurate when the packet length is extremely short, and/or the tolerable bit error rate is extremely low. For further discussions regarding the accuracy of the NA, please refer to [24, 25, 26]., which is given by [27, Theorem 78]
@@
This paper [15] [30] [16, 17, 18, 19, 20, 31] [32, 33]
Multi-user $\surd$ $\surd$ $\surd$ $\surd$ -
Ch. disp. in [34] $\surd$ $\surd$ $\surd$ - -
EE $\surd$ $\surd$ - - -
MIMO $\surd$ - - - -
Multiple streams $\surd$ - - - -
STAR-RIS $\surd$ - - - -
Transmission delay $\surd$ - - - -
@@ I-C Contribution
To elaborate, our optimization framework is flexible and can be utilized in a wide range of MU-MIMO URLLC systems aided by RISs. Additionally, the framework may be used for solving a broad spectrum of optimization problems for which the objective function and/or constraints can be, but are not limited to, linear functions of the rates and/or EE of users. The convergence of our framework is ensured towards a stationary point for the general optimization problem, when the feasibility set for the RIS elements conforms to a convex set. We consider a multi-cell MIMO BC as an example of the networks that our framework can be applied to. Furthermore, we consider both the EE and SE metrics as well as the transmission delay for investigating the performance of RIS in MU-MIMO URLLC systems. For the SE metric, we consider the sum rate and the minimum rate of the users, which are among the most common performance metrics for SE. Moreover, we evaluate the EE by optimizing the GEE and the minimum EE of users. Note that the sum rate and global EE are pivotal overall system performance metrics. By contrast, the minimum rate and the EE of specific users consider the individual performance of the users and can provide reasonable rate/EE-fairness among the users since typically all the users are allocated similar rate/EE when the minimum rate/EE is maximized. Thus, considering all these metrics can provide a complete picture of the performance of MU-MIMO URLLC systems aided by RISs. Moreover, we make realistic assumptions regarding the channel models and the feasible sets of the RIS coefficients for appropriately examining the RIS performance.
@@ II System model
Our proposed framework can be applied to a large family of RIS-assisted MU-MIMO URLLC systems that treat interference as noise at the receivers. As an example of such MU-MIMO systems, we consider a multi-cell MIMO RIS-assisted downlink (DL) BC comprising $L$ BSs, as shown in Fig. 1. We assume that BS $l$ has $N_{BS,l}$ DL transmission antennas (TAs) and serves $K_{l}$ multiple-antenna users. The $k$-th user associated to BS $l$, denoted as Ulk, has $N_{u,lk}$ receive antennas (RAs). Additionally, we assume that there are $M$ reflective passive RISs to assist the BSs, and the $m$-th RIS has $N_{RIS,m}$ elements. Furthermore, we assume perfect, instantaneous, and global CSI, consistent with many other studies on RISs [5, 6, 8, 35, 28, 36]. This assumption is also commonly used in the development of resource allocation solutions for URLLC systems [37, 38, 31, 39, 40, 14]. These solutions are particularly applicable in systems with large channel coherence time, where the channel state remains stable for extended periods, making channel estimation easier and more accurate. In such systems, resource allocation solutions can be reused across multiple time slots, and the pilot overhead required for acquiring CSI and computing the solutions is relatively low. Additionally, investigating the performance of RISs under the assumption of perfect CSI helps to illustrate the essential tradeoffs in system design and provides an upper bound of the system performance.
Figure 1: A multi-cell BC assisted by RISs.
@@ II-A RIS Model
We consider two nearly-passive RIS architectures, namely reflective RIS and STAR-RIS, and employ the RIS model of [35] for the MIMO multi-cell BC.
II-A1 Reflective RIS
The channel matrix between BS $i$ and Ulk as a function of the RIS matrices is given by
$\mathbf{H}_{lk,i}\!\left(\{\bf{\Psi}\}\right)\!\!=\!\!\!\underbrace{\sum_{m=1}^{M}\!\!\mathbf{G}_{lk,m}{\bf\Psi}_{m}\mathbf{G}_{m,i}}_{\text{Link through RIS}}+\!\!\!\!\underbrace{\mathbf{F}_{lk,i}}_{\text{Direct link}}\!\!\!\!\in\mathbb{C}^{N_{u,lk}\times N_{BS,l}}\!,$ (2)
where $\mathbf{F}_{lk,i}\in\mathbb{C}^{N_{u,lk}\times N_{BS,i}}$ is the channel matrix between the $i$-th BS and Ulk, $\mathbf{G}_{lk,m}\in\mathbb{C}^{N_{u,lk}\times N_{RIS,m}}$ is the channel matrix between the $m$-th RIS and Ulk, and $\mathbf{G}_{m,i}\in\mathbb{C}^{N_{RIS,m}\times N_{BS,i}}$ is the channel matrix between the $i$-th BS and the $m$-th RIS. Additionally, $\{{\bf\Psi}\}=\{{\bf\Psi}_{m}\}_{m=1}^{M}$ denotes the set of all coefficients of RISs, where ${\bf\Psi}_{m}\in\mathbb{C}^{N_{RIS,m}\times N_{RIS,m}}$ is a diagonal matrix, containing the vector of reflecting coefficients of the $m$-th RIS
${\bf\Psi}_{m}=\text{diag}\left(\psi_{m_{1}},\psi_{m_{2}},\cdots,\psi_{m_{N_{RIS,m}}}\right).$
Assuming having nearly passive RISs, the absolute value of the RIS coefficients cannot be greater than 1, which results in the following set for the feasible RIS coefficients [3, Eq. (11)]
@@ II-A1 Reflective RIS
In this feasibility set, the amplitude and phase of each RIS element are assumed to be independent optimization variables, which might not be realistic. Another common assumption is that the RIS coefficients have to adhere to the unit modulus constraint [4, 3, 6, 7, 41, 35, 42], which leads to
@@
II-A2 STAR-RIS
STAR-RIS provides an omni-directional $360^{\circ}$ full-place coverage. In STAR-RIS, each component can operate in both reflection and transmission mode [43, 44]. Thus, there are two complex-valued optimization parameters per element, when STAR-RIS is employed. We denote the reflection/transmission coefficient for the $n$-th element of the $m$-th RIS as $\psi_{m_{n}}^{r}/\psi_{m_{n}}^{t}$. Based on the position of the user with respect to STAR-RIS, the STAR-RIS can optimize the channel of the user only through the reflection or transmission coefficients. Therefore, the channel between BS $i$ and Ulk is
@@ II-A2 STAR-RIS
where we have ${\bf\Psi}_{m}^{r}=\text{diag}\left(\psi_{m_{1}}^{r},\psi_{m_{2}}^{r},\cdots,\psi_{m_{N_{RIS,m}}^{r}}\right)$ and ${\bf\Psi}_{m}^{t}=\text{diag}\left(\psi_{m_{1}}^{t},\psi_{m_{2}}^{t},\cdots,\psi_{m_{N_{RIS,m}}^{t}}\right)$. Assuming operating in a passive mode, the absolute values of the reflection and transmission coefficients have to satisfy
$|\psi_{m_{n}}^{r}|^{2}+|\psi_{m_{n}}^{t}|^{2}\leq 1,\hskip 28.45274pt\forall m,n,$ (6)
@@
There are three different STAR-RIS schemes, including the energy splitting (ES), mode switching (MS) and time switching (TS) schemes [45, 44]. Since the main focus of this work is on evaluating the impact of employing multiple streams per user regimes, we consider only the MS scheme. Note that the MS scheme has a lower implementation complexity than the ES scheme, but its performance may be comparable to that of the ES scheme as shown in, e.g., [46, 47, 48]. The framework proposed in this treatise can be extended to include the ES and TS schemes by following an approach similar to [47, 48].
II-A3 Brief comparison of reflective RIS and STAR-RIS
The main difference between a reflective RIS and a STAR-RIS is in the coverage area of these RIS architectures, which makes each suitable for a different set of applications. STAR-RIS can provide omni-directional coverage, while reflective RIS can assist the communication between a BS and a user only if they are in the reflection space of the RIS. Thus, when the RIS can be positioned for ensuring that all the transceivers are in the reflection space of the RIS, a reflective RIS could be a more suitable option. However, when the BS is located outdoors and supports both indoor and outdoor users, STAR-RISs are preferable.
From an optimization point of view, each STAR-RIS element has two complex-valued coefficients. Thus, if the ES scheme is employed, algorithms conceived for STAR-RIS could have slightly higher computational complexities compared to reflective RIS. However, if MS or TS schemes are utilized, each RIS element operates either in the reflection mode or in the transmission mode. Thus, only one coefficient for each STAR-RIS element is optimized, which reduces the computational complexity to the same order as that of reflective RIS algorithms.
The channel matrices are assumed to be linear/affine functions of the RIS elements in both reflective RIS and STAR-RIS. To simplify the notations/equations, we remove this dependency and subsequently denote the channels as $\mathbf{H}_{lk,i}$ for all $l,k,i$, hereafter. Additionally, we denote the set of the feasible RIS elements as $\mathcal{T}$, unless we refer to a specific set. Note that there are other RIS technologies and/or more practical feasibility sets for STAR/reflective RIS as mentioned in [3, 49], which should be considered in future studies.